OpenStudy (anonymous):

square root of (9+4x) x square root of (9+4x) =?

OpenStudy (anonymous):

Is that "x" in the middle a symbol for multiplication? Like this? : \[\sqrt{9+4x} * \sqrt{9+4x}\]

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

I was tempted to foil.., but then I think I also am tempted to make it 9 + 4x.

OpenStudy (anonymous):

When you are multiplying the square root of the exact same thing, the square root disappears. You are multiplying two exact same quantities together, so something times the same something = (something) squared.\[\sqrt{9+4x} * \sqrt{9+4x} = (\sqrt{9+4x})^{2}\] If you have a square root to the power of 2, that power of 2 makes the square root go away. So you are just left with 9+4x.

OpenStudy (anonymous):

Ah alright thanks!!! :)

OpenStudy (anonymous):

Do you know why a power of 2 makes a square root go away? Because a square root just means "the power of 1/2." So a square root of something means that something to the power of 1/2. \[\sqrt{9+4x}=(9+4x)^{1/2}\] Now to square that, we have:\[((9+4x)^{1/2})^{2}\] Multiplying those exponents gives 1/2 * 2, which is just equal to 1. So (9+4x) to the power of 1 is just itself, (9+4x). Just a side note ;) You're welcome!

OpenStudy (anonymous):

Ah, this makes a lot of sense. Thanks so much!